OpenStudy (anonymous):

Is (-3,4) (0,19) and (2,9) ordered pair of the solution y=-5x+19

OpenStudy (shinalcantara):

For to find out which coordinates/points are in the line, all you have to do is plug in the values of x and see if the value of y will coincide with that of its ordinate. Let's take the first one y=-5x + 19 (-3,4) Plug in th evalue of x y = -5(-3) + 19 y = ?

OpenStudy (shinalcantara):

@AndreaGreer

OpenStudy (anonymous):

So yes an ordered pair (-3,4)?

OpenStudy (shinalcantara):

what's the value of y from above equation?

OpenStudy (anonymous):

It don't tell me it just shows y=-5x+19 then they want to know if (-3,4) is an ordered pair

OpenStudy (shinalcantara):

that's why were getting the value of y when x is plugged in to the equation to know whether the coordinates are the points of the line or not

OpenStudy (anonymous):

Yes so yes (-3,4) would be ordered pair

OpenStudy (shinalcantara):

Taking the first coordinates you'll have y=-5x + 19 (-3,4) Plug in th evalue of x y = -5(-3) + 19 y = 34 ----not ok! since 4 is the y-coordinate For the second point (0,19) y=-5x + 19 y = -5(0) + 19 y = 19 ok! since 19 is the ordinate of the point For the third one (2,9) y=-5x + 19 y = -5(2) + 19 y = -10 + 19 y = 9 ---ok! since in the ordered pair (2,9), 9 is the y-coordinate